# Runtime Safety through Adaptive Shielding: From Hidden Parameter Inference to Provable Guarantees

### v1

@@ Appendix E Ablation Study on Safety Bonus and Sampling Size
We evaluate the hyperparameter sensitivity of our method, focusing on the safety bonus and sampling size.
To assess the safety bonus, we fix the sampling size at 10 and vary the safety bonus in ${0.05,0.1,0.5,1}$. Figures 6 and 7 demonstrate that a higher safety bonus reduces cost violations, consistent with the safety-regularized objective, as it increasingly incentivizes policies to minimize safety violations.
Figure 6: Ablation study on safety bonus (Car Robot, TRPO-Lag Base): (Top) Average episodic return over training steps, with curves for safety bonus values {0.05, 0.1, 0.5, 1}. (Bottom) Violation rate over training steps, decreasing with higher safety bonuses (0.5, 1). Sampling number fixed at 10.
Figure 7: Ablation study on safety bonus (Point Robot, TRPO-Lag Base): (Top) Average episodic return over training steps, with curves for safety bonus values {0.05, 0.1, 0.5, 1}. (Bottom) Violation rate over training steps, decreasing with higher safety bonuses (0.5, 1). Sampling number fixed at 10.
Figure 8: Ablation study on sampling size (Car Robot, TRPO-Lag Base): (Top) Average episodic return over training steps, with curves for sampling numbers {5, 10, 50, 100}. (Bottom) Violation rate over training steps, stable at low levels across sampling numbers. Safety bonus fixed at 1.
Figure 9: Ablation study on sampling size (Point Robot, TRPO-Lag Base): (Top) Average episodic return over training steps, with curves for sampling numbers {5, 10, 50, 100}. (Bottom) Violation rate over training steps, stable at low levels across sampling numbers. Safety bonus fixed at 1.
To investigate the impact of sampling size during training, we fix the safety bonus at 1 and vary the sampling size in ${5,10,50,100}$. We observe that sampling size influences the early training phase, where an untrained policy benefits from larger samples to explore diverse actions. Figures 8 and 9 demonstrate that, as training advances, policies with varying sampling sizes converge to comparable performance in both reward and cost. These results support the robustness of our method.
Figure 10: Ablation study on sampling size in OOD environments (PPO-Lag Base): $x$-axis denotes cost, $y$-axis denotes reward. (Top) Car Robot. (Bottom) Point Robot.
Figure 11: Ablation study on sampling size in OOD environments (TRPO-Lag Base): $x$-axis denotes cost, $y$-axis denotes reward. (Top) Car Robot. (Bottom) Point Robot.
We evaluate the impact of sampling size during testing, varying it across ${5,10,50,100}$. Figures 10 and 11 show that sampling size affects performance. For instance, the Car robot exhibits performance variation with sampling size, whereas the Point robot is relatively robust to changes in sampling size. However, the figures demonstrate that our method, with different sampling sizes, remains on the Pareto frontier.
Appendix F Ablation on Representation
We investigate the function encoder’s representation of varying underlying dynamics $T_{\theta}$. We evaluate two representations for handling hidden parameters in our safe RL framework: the Oracle representation, where the hidden parameter $\theta$ (a scaling factor for environmental dynamics such as density and damping) is directly provided to the policy by concatenating it with the state input, and the Function Encoder (FE) representation, which uses a function encoder $\hat{f}_{\text{FE}}$ to infer the underlying dynamics $T_{\theta}$, with coefficients of pretrained basis functions serving as the representation. These representations are tested to assess the function encoder’s ability to adapt to varying dynamics in Safety Gymnasium tasks.
Figures 12 and 13 show that the function encoder’s representation is often comparable to the oracle representation and, in some cases, outperforms it. For instance, in the Car-Goal task, the function encoder’s representation surpasses the oracle representation.
The function encoder leverages neural basis functions to represent the space of varying dynamics $\left\{T_{\theta}\right\}_{\theta\in\Theta}$. For instance, just as the $\mathbb{R}^{2}$ plane is spanned by linear combinations of basis vectors $(0,1)$ and $(1,0)$, the dynamics space is captured by neural basis functions, making their coefficients highly informative. This representation often transitions smoothly, as shown in [Ingebrand et al., 2024b], promoting policy effective adaptation to dynamic changes. Consequently, our function encoder’s representation frequently matches or surpasses oracle representation performance.
@@ Appendix F Ablation on Representation
Figure 12: Ablation study on Representation (Car robot). "Oracle-" refers to a policy directly informed of hidden parameters, while "FE-" denotes the function encoder’s representation derived from observations.
Figure 13: Ablation study on Representation (Point robot). "Oracle-" refers to a policy directly informed of hidden parameters, while "FE-" denotes the function encoder’s representation derived from observations.
Appendix G Why Not $Q_{C}$, But $Q_{\text{safe}}$?

### v2

@@ Appendix E Ablation Study on Safety Bonus and Sampling Size
We evaluate the hyperparameter sensitivity of our method combined with RCPO, focusing on safety bonus $\alpha$ and sampling size $s$.
Varying Safety Bonus $\alpha$ with RCPO. To assess the sensitivity of the safety bonus, we vary the safety bonus $\alpha$ across $\{0.05,0.1,0.5,1.0\}$. We observe that performance with SRO often improves both reward and safety. This is mainly because SRO encourages the policy to select safe actions while also exploring under-explored actions. However, SRO does not consistently enhance both reward and safety; instead, it frequently improves either reward or cost. These findings align with the theoretical results presented in Appendix B.
Algo. Env. Point-Goal Point-Button Point-Push Point-Circle
R $\uparrow$ C(%) $\downarrow$ R $\uparrow$ C(%) $\downarrow$ R $\uparrow$ C(%) $\downarrow$ R $\uparrow$ C(%) $\downarrow$
RCPO 16.35±1.14 2.50±0.02 8.97±0.56 1.69±0.23 0.12±0.15 0.63±0.28 30.78±5.55 2.47±0.51
SRO + RCPO ($\alpha$=0.05) 17.49±1.82 2.89±0.52 10.38±1.89 1.76±0.24 0.33±0.16 0.88±0.48 24.33±4.58 1.55±0.03
SRO + RCPO ($\alpha$=0.1) 17.56±1.35 2.63±0.25 10.28±2.00 1.64±0.23 0.32±0.15 0.65±0.37 25.95±4.37 0.98±0.02
SRO + RCPO ($\alpha$=0.5) 17.58±0.56 2.49±0.34 11.26±1.28 1.82±0.31 0.33±0.12 0.61±0.10 25.18±4.50 1.28±0.29
SRO + RCPO ($\alpha$=1.0) 15.93±0.78 2.37±0.11 9.05±1.43 1.77±0.43 0.33±0.04 0.82±0.56 24.97±3.74 1.55±0.34
Algo. Env. Car-Goal Car-Button Car-Push Car-Circle
R $\uparrow$ C(%) $\downarrow$ R $\uparrow$ C(%) $\downarrow$ R $\uparrow$ C(%) $\downarrow$ R $\uparrow$ C(%) $\downarrow$
RCPO 15.64±2.27 2.23±0.33 5.74±0.34 1.89±0.22 -0.09±0.13 0.60±0.09 11.83±0.49 1.58±0.57
SRO + RCPO ($\alpha$=0.05) 18.54±0.65 2.13±0.14 7.01±0.51 1.80±0.26 -0.10±0.17 0.88±0.25 11.29±0.28 1.29±0.19
SRO + RCPO ($\alpha$=0.1) 15.88±1.12 2.31±0.66 7.68±1.21 1.85±0.19 0.06±0.09 0.68±0.41 11.47±0.53 1.51±0.61
SRO + RCPO ($\alpha$=0.5) 17.33±0.58 2.55±0.50 6.71±1.11 1.88±0.46 -0.05±0.12 0.59±0.39 11.19±0.25 1.52±0.47
SRO + RCPO ($\alpha$=1.0) 16.73±2.46 1.73±0.29 7.89±0.95 1.95±0.47 -0.14±0.18 0.67±0.19 11.50±0.16 2.15±0.50
Table 2: Ablation Study on the Varying Effects of Safety Bonus $\alpha$ on Safety and Performance. Best performances (highest return and lowest cost rate) are highlighted in bold.
Varying Sampling Numbers $s$ with RCPO. We evaluate the impact of sampling size, varying it across $\{5,10,20,50\}$ by fixing safety bonus $\alpha=1.0$. Table 3 demonstrates that sampling size influences performance. Sampling numbers exhibit no consistent pattern due to randomness in the sampling procedure. This arises primarily from high prediction errors, which often lead to incorrect action sampling, even within conformal prediction boundaries. For example, a large error widens the conformal interval range, causing the shield to include numerous sampled actions to meet the probabilistic guarantee. However, even when selecting actions based on safety scores, these high errors may inaccurately represent safe actions.
Algo. Env. Point-Goal Point-Button Point-Push Point-Circle
R $\uparrow$ C(%) $\downarrow$ R $\uparrow$ C(%) $\downarrow$ R $\uparrow$ C(%) $\downarrow$ R $\uparrow$ C(%) $\downarrow$
RCPO 16.35±1.14 2.50±0.02 8.97±0.56 1.69±0.23 0.12±0.15 0.63±0.28 30.78±5.55 2.47±0.51
Shield + SRO (s=5) 12.83±1.91 2.27±0.60 8.68±2.51 2.43±0.39 0.13±0.27 0.51±0.27 26.76±4.71 1.38±0.45
Shield + SRO (s=10) 16.40±0.73 2.29±0.21 8.22±3.31 1.98±0.72 0.36±0.40 0.65±0.04 29.97±0.43 2.16±0.43
Shield + SRO (s=20) 14.74±1.18 1.96±0.37 7.43±1.55 2.19±0.45 0.03±0.18 0.90±0.53 26.76±5.40 1.81±0.36
Shield + SRO (s=50) 14.37±0.65 2.34±0.23 7.34±1.27 1.58±0.47 0.31±0.20 0.53±0.37 28.24±4.54 1.87±0.63
Algo. Env. Car-Goal Car-Button Car-Push Car-Circle
R $\uparrow$ C(%) $\downarrow$ R $\uparrow$ C(%) $\downarrow$ R $\uparrow$ C(%) $\downarrow$ R $\uparrow$ C(%) $\downarrow$
RCPO 15.64±2.27 2.23±0.33 5.74±0.34 1.89±0.22 -0.09±0.13 0.60±0.09 11.83±0.49 1.58±0.57
Shield + SRO (s=5) 14.43±0.27 2.12±0.15 5.24±0.90 1.86±0.28 0.09±0.07 0.75±0.24 11.55±0.29 1.28±1.02
Shield + SRO (s=10) 13.65±0.35 1.84±0.14 5.52±0.90 1.92±0.43 0.23±0.09 0.86±0.24 11.61±0.67 1.44±0.84
Shield + SRO (s=20) 15.69±0.47 2.51±0.50 4.93±0.81 1.99±0.20 0.04±0.18 0.61±0.41 11.44±0.20 1.88±1.02
Shield + SRO (s=50) 13.23±1.73 2.65±0.86 4.88±1.37 1.52±0.17 0.00±0.10 0.50±0.20 11.08±0.37 1.49±0.74
Table 3: Ablation Study on the Varying Effects of Sampling Numbers $s$ on Safety and Performance with fixed safety bonus $\alpha=1.0$. Best performances (highest return and lowest cost rate) are highlighted in bold.
Appendix F Adaptive Shielding and Safety-Regularized Optimization with PPO-Lag
Our method is compatible with a wide range of RL algorithms, as it wraps the policy with a shielding layer and incorporates an augmented term based on $Q_{C}$ and $V_{C}$, which are commonly used in safe RL algorithms. To demonstrate this, we examine its impact when applied to a PPO-Lagrangian-based policy. Here, we study PPO-Lag augmented with SRO and Adaptive Shielding mechanisms. The baseline PPO-Lag method is provided with access to the hidden parameter $\phi$. First, we analyze how SRO affects PPO-Lagrangian method. Then, we show how Shielding mechanism combined with SRO affects PPO-Lagrangian method. All results shown represent the mean reward and cost rate over the last 20 epochs of training across seeds.
Safety-Regularized Optimization with PPO-Lag. Table 4 demonstrate that SRO generally enhances safety. In Point Robot case, a clear pattern shows that higher safety bonus values $\alpha$ improve safety the most, while lower $\alpha$ have minor effects. Moreover, adding SRO does not degrade reward performance substantially, with cost violations improving by up to 20% (Point-Button) to as much as 520% (Point-Circle). Meanwhile, reward degradation occurs only in Point-Circle and Car-Button environments; in many other cases, SRO improves not only safety but also the reward signal. This is primarily because our augmented objective is bounded in (1,0](-1,0], mildly influencing the training objective to compensate for actions leading to zero long-term cost violations and under-explored actions, without causing significant shifts during training due to the bounded values.
Algo. Env. Point-Goal Point-Button Point-Push Point-Circle
R $\uparrow$ C(%) $\downarrow$ R $\uparrow$ C(%) $\downarrow$ R $\uparrow$ C(%) $\downarrow$ R $\uparrow$ C(%) $\downarrow$
PPOLag 18.20±0.78 3.35±0.20 10.83±1.56 2.94±0.22 0.24±0.19 1.15±0.27 30.90±3.52 13.38±6.74
SRO + PPOLag ($\alpha$=0.05) 19.41±0.45 2.97±0.25 10.92±1.65 3.16±0.54 0.19±0.08 1.16±0.10 25.20±2.31 4.56±1.87
SRO + PPOLag ($\alpha$=0.1) 17.65±0.26 2.52±0.07 11.63±0.80 3.04±0.58 0.53±0.27 1.32±0.36 27.79±1.04 2.93±1.49
SRO + PPOLag ($\alpha$=0.5) 17.12±0.97 2.51±0.50 10.49±0.68 2.88±0.32 0.32±0.03 0.85±0.24 26.50±3.41 3.58±1.07
SRO + PPOLag ($\alpha$=1.0) 16.98±1.36 2.79±0.36 9.69±0.55 2.44±0.35 0.36±0.20 0.59±0.16 26.07±4.88 2.15±1.27
Algo. Env. Car-Goal Car-Button Car-Push Car-Circle
R $\uparrow$ C(%) $\downarrow$ R $\uparrow$ C(%) $\downarrow$ R $\uparrow$ C(%) $\downarrow$ R $\uparrow$ C(%) $\downarrow$
PPOLag 15.85±1.01 3.36±0.46 8.62±1.26 3.58±0.36 0.02±0.09 1.14±0.12 11.48±0.38 2.57±0.23
SRO + PPOLag ($\alpha$=0.05) 16.10±1.07 3.08±0.11 8.37±1.69 3.31±1.13 0.02±0.07 0.95±0.20 11.24±0.13 2.32±0.36
SRO + PPOLag ($\alpha$=0.1) 16.81±1.17 3.16±0.18 8.53±1.64 3.56±0.36 0.02±0.04 1.07±0.09 11.46±0.45 2.00±0.20
SRO + PPOLag ($\alpha$=0.5) 15.50±1.11 3.18±0.34 8.24±0.54 3.14±0.43 -0.02±0.10 1.14±0.29 11.52±0.47 2.02±0.88
SRO + PPOLag ($\alpha$=1.0) 16.41±0.57 3.60±0.53 8.11±1.17 3.78±0.16 -0.01±0.10 0.99±0.23 11.42±0.56 1.74±0.60
Table 4: Ablation Study on the Varying Effects of Safety Bonus $\alpha$ on Safety and Performance. Best performances (highest return and lowest cost rate) are highlighted in bold.
Adaptive Shielding with PPO-Lag. For sampling numbers, unlike the safety bonus $\alpha$, no consistent pattern emerges due to randomness in the sampling procedure. The same reasoning outlined in Appendix E applies here. Thus, reducing prediction errors and mitigating the inherent randomness in the sampling process represent key areas for future research to enhance shielding-based approaches.
Algo. Env. Point-Goal Point-Button Point-Push Point-Circle
R $\uparrow$ C(%) $\downarrow$ R $\uparrow$ C(%) $\downarrow$ R $\uparrow$ C(%) $\downarrow$ R $\uparrow$ C(%) $\downarrow$
PPOLag 18.20±0.78 3.35±0.20 10.83±1.56 2.94±0.22 0.24±0.19 1.15±0.27 30.90±3.52 13.38±6.74
Shield + SRO (s=5) 17.41±1.96 2.84±0.36 10.27±1.30 2.87±0.36 0.38±0.11 1.01±0.12 25.74±3.26 4.25±1.30
Shield + SRO (s=10) 18.38±0.84 2.86±0.25 10.18±0.86 2.87±0.23 0.20±0.06 0.92±0.36 27.34±8.09 2.42±2.48
Shield + SRO (s=20) 17.10±2.28 2.75±0.32 8.84±0.05 2.67±0.27 0.15±0.12 0.82±0.10 28.83±3.36 2.03±0.87
Shield + SRO (s=50) 17.44±0.90 2.60±0.26 11.23±3.43 3.12±1.07 0.37±0.10 1.27±0.64 27.55±4.32 2.43±1.54
Algo. Env. Car-Goal Car-Button Car-Push Car-Circle
R $\uparrow$ C(%) $\downarrow$ R $\uparrow$ C(%) $\downarrow$ R $\uparrow$ C(%) $\downarrow$ R $\uparrow$ C(%) $\downarrow$
PPOLag 15.85±1.01 3.36±0.46 8.62±1.26 3.58±0.36 0.02±0.09 1.14±0.12 11.48±0.38 2.57±0.23
Shield + SRO (s=5) 15.56±0.47 3.66±0.64 7.57±0.39 3.28±0.12 0.03±0.11 1.48±0.46 10.38±0.39 1.43±0.70
Shield + SRO (s=10) 16.87±5.59 3.79±0.24 7.62±0.91 3.35±0.57 0.05±0.10 1.12±0.30 10.15±0.67 1.01±0.27
Shield + SRO (s=20) 15.97±1.36 3.43±0.83 7.56±1.01 2.64±0.47 0.09±0.14 1.22±0.36 10.33±0.15 0.93±0.09
Shield + SRO (s=50) 13.93±2.09 3.48±0.50 6.25±1.47 3.22±0.41 0.05±0.10 1.01±0.55 11.38±0.54 1.49±0.07
Table 5: Ablation Study on the Varying Effects of Sampling Numbers $s$ on Safety and Performance with fixed safety bonus $\alpha=1.0$. Best performances (highest return and lowest cost rate) are highlighted in bold.
@@ Appendix F Adaptive Shielding and Safety-Regularized Optimization with PPO-Lag
Figure 3: Ablation study on Representation. "Oracle-" refers to a policy directly informed of hidden parameters, while "FE-" denotes the function encoder’s representation derived from observations.
Appendix G Function Encoders and Its Representation
For completeness, we present the necessary background on function encoders. For more information, see (ingebrand2024zeroshot_ode; ingebrand2025functionencodersprincipledapproach).
Overview of Function Encoders. Function encoders, as introduced by (ingebrand2025functionencodersprincipledapproach), provide a principled framework for representing tasks in a Hilbert space $\mathcal{H}$ through a finite set of neural network-based basis functions $\{g_{1},\ldots,g_{k}\}$, each parameterized by $\theta_{j}$. A task function $f\in\mathcal{H}$ is approximated as a linear combination of these basis functions:
$f(x)=\sum_{j=1}^{k}b_{j}g_{j}(x\mid\theta_{j}),$
where $b_{j}$ are coefficients tailored to the specific task. This approach enables efficient representation of complex functions, such as those encountered in reinforcement learning or classification, by learning a versatile basis that spans the function space. By defining appropriate inner products, function encoders can generalize to various function spaces, including probability distributions for classification tasks. During supervised training, the neural basis functions $\{g_{j}\}$ are optimized such that $\sum_{j}b_{j}g_{j}(x_{i})\approx f(x_{i})$ for all training points $(x_{i},f(x_{i}))$. As expected from functional analysis, using more basis functions increases expressive power and allows modeling more complex dynamics. Empirically, we demonstrate that a larger number of basis functions yields richer representations showing faster convergence of the dynamics prediction loss. Models with fewer basis functions eventually reach similar final performance but require more training epochs (Appendix H).
The training process consists of two phases: offline training of the basis functions and online inference to compute task-specific coefficients. The offline phase optimizes the basis to minimize reconstruction error across a set of source datasets, while the online phase efficiently computes coefficients for new tasks using the learned basis.
Training Function Encoders via Least Squares. Function encoder is trained by using a least-squares optimization approach (ingebrand2025functionencodersprincipledapproach). Given a set of task functions $\{f_{1},\ldots,f_{n}\}$, the goal is to learn a set of basis functions $\{g_{1},\ldots,g_{k}\}$ parameterized by $\theta$ and these basis functions represent the task functions with varying coefficients $b$. In our implementation, each task function $f_{\ell}$ is defined by fixing one set of hidden parameters: gravity, mass, damping, density, friction.
The training procedure iteratively minimizes a loss function comprising two components: a reconstruction loss and a regularization term. For each task function $f_{\ell}$, we compute coefficients $b^{\ell}=[b_{1}^{\ell},\ldots,b_{k}^{\ell}]^{T}$ that best approximate the target function $f_{\ell}$ as:
$b^{\ell}=\left[\begin{array}[]{ccc}\langle g_{1},g_{1}\rangle_{\mathcal{H}}&\cdots&\langle g_{1},g_{k}\rangle_{\mathcal{H}}\\ \vdots&\ddots&\vdots\\ \langle g_{k},g_{1}\rangle_{\mathcal{H}}&\cdots&\langle g_{k},g_{k}\rangle_{\mathcal{H}}\end{array}\right]^{-1}\left[\begin{array}[]{c}\langle f_{\ell},g_{1}\rangle_{\mathcal{H}}\\ \vdots\\ \langle f_{\ell},g_{k}\rangle_{\mathcal{H}}\end{array}\right],$
where $\langle\cdot,\cdot\rangle_{\mathcal{H}}$ denotes the inner product in the Hilbert space, estimated via Monte Carlo integration over collected data points $\{(x_{1},f_{\ell}(x_{1})),((x_{2},f_{\ell}(x_{2})),\cdots,((x_{N},f_{\ell}(x_{N}))\}$. The reconstructed function is then $\hat{f}_{\ell}=\sum_{j=1}^{k}b_{j}^{\ell}g_{j}$. The reconstruction loss is defined as:
$L=\frac{1}{n}\sum_{i=1}^{n}\|f_{i}-\hat{f}_{i}\|_{\mathcal{H}}^{2},$
which measures the average squared error between the true and approximated functions. To ensure the basis functions remain well-conditioned, a regularization term is added to the loss function:
$L_{\text{reg}}=\sum_{i=1}^{k}\left(\|g_{i}\|_{\mathcal{H}}^{2}-1\right)^{2},$
which encourages the basis functions to have unit norm. For a learning rate $\alpha$, the parameters $\theta$ are updated via gradient descent: $\theta\leftarrow\theta-\alpha\nabla_{\theta}(L+L_{\text{reg}})$, until convergence.
Empirical Evaluation of the Representation. We investigate the function encoder’s representation of varying underlying dynamics $T_{\phi}$. We evaluate two representations for handling hidden parameters in our safe RL framework: Oracle representation, where the hidden parameter $\phi$ (a scaling factor for environmental dynamics such as density and damping) is directly provided to the policy by concatenating it with the state input, and Function Encoder (FE) representation, which uses a function encoder $\hat{f}_{\text{FE}}$ to infer the underlying dynamics $T_{\phi}$, with coefficients of pretrained basis functions serving as the representation. These representations are tested to assess the function encoder’s ability to adapt to varying dynamics in Safety Gymnasium tasks. Regarding training hyperparameters, we employed three number of basis functions for the function encoder, trained over 1000 epochs on a dataset of 1000 episodes. Batch size was set to 256. Figure 3 show that the function encoder’s representation is often comparable to the oracle representation and, in some cases, outperforms it. The function encoder leverages neural basis functions to represent the space of varying dynamics $\left\{T_{\phi}\right\}_{\phi\in\Phi}$. For instance, just as the $\mathbb{R}^{2}$ plane is spanned by linear combinations of basis vectors $(0,1)$ and $(1,0)$, the dynamics space is captured by neural basis functions, making their coefficients highly informative. This representation often transitions smoothly, as shown in (tyler2024zero_shot), promoting policy effective adaptation to dynamic changes. Consequently, our function encoder’s representation frequently matches or surpasses oracle representation performance.
Figure 4: (a) Illustration of how function encoders obtain proxy representations of the underlying hidden parameters using online samples. (b) A naive approach using transformer encoder to infer hidden parameters from a sequence of online samples. MLP stands for multi-layer perceptron in the Figure.
Appendix H Shielding with Alternative Dynamics Predictors
In this section, we present ablation studies on different dynamics predictors by replacing the function encoder with alternative prediction models. We consider three predictors: a naive transformer-based dynamics model, a probabilistic ensemble model (PEM), and a multilayer perceptron (MLP). Since PEM and MLP are not designed to infer hidden environment parameters $\phi$ from context alone, we provide them with the true $\phi$ as additional input. Following prior baselines that assume access to the environment parameter $\phi$, we refer to these models as Oracle-PEM and Oracle-MLP.
As another baseline described in Figure 4, we use a transformer encoder to process the current episode’s trajectory
$\tau_{n}=\left(s_{0},a_{0},s_{1},a_{1},\ldots,s_{n}\right)$
and extract a latent representation $b_{\phi}=\operatorname{Enc}\left(\tau_{n}\right)$ as a proxy for the hidden parameters $\phi$. For a fair comparison to function encoder, the transformer encoder is given $100$ samples to infer $b_{\phi}$, which is then concatenated to the state $s_{t}$ and fed into the predictor to generate $\hat{s}_{t}=(s_{t},b_{\phi})$. All dynamics predictor are designed to hold a comparable parameter numbers (270k-280k). All dynamics predictors are trained on $1,000$ in-distribution episodes and evaluated on $200$ out-of-distribution test episodes, which are not used during training.
We first report next state prediction performance on evaluation OOD dataset across all dynamics predictors. We then compare the naive transformer combined with Shield and Shield + SRO against the function encoder combined with the same shielding mechanisms. Note that SRO alone does not use dynamics prediction at deployment.
Figure 5 shows that the function encoder’s next-state prediction accuracy closely matches that of Oracle-PEM and Oracle-MLP. This observation is consistent with prior findings (tyler2024zero_shot; ingebrand2024zeroshot_ode), which also report that function-encoder-based models can approximate the performance of oracle-informed predictors.
Instead of inferring the hidden parameters $\phi$ through neural basis functions, the Transformer encodes the observation sequence $\tau=(s_{0},a_{0},s_{1},\ldots,s_{n})$ and projects the final embedding into a low-dimensional parameter estimate $b_{\phi}$. Since the true hidden parameters $\phi$’s dimension in safe-navigation domain is 4 (damping, mass, inertia, friction), we evaluate projection dimensions of $3,6,9$, and $12$.
Figure 6 shows an interesting pattern. For the function encoder, increasing the number of basis functions shows early-epoch convergence because the representation has higher capacity, but all configurations eventually converge to similar accuracy, which aligns with the supervised nature of the objective and dataset limitations. In contrast, the Transformer-based encoder does not exhibit a consistent relationship between projection dimension and prediction quality, and its overall performance is less stable.
Finally, Figure 7 shows that the function encoder consistently yields superior performance, achieving higher rewards and fewer constraint violations. This improvement originates from the combination of $(i)$ a more stable and expressive learned representation and $(ii)$ higher next-state prediction accuracy, both of which enhance shielding framework.
Figure 5: Ablation study evaluating the performance of various dynamics predictors in forecasting the next state. The $y$-axis denotes the average per-sample $\ell_{1}$-norm error between the true and predicted next states on the test dataset.
Figure 6: We evaluate how the dimension of the inferred representation $b_{\phi}$, which is a proxy to hidden parameters $\phi$. We vary the dimension across $\{3,6,9,12\}$, motivated by the ground-truth hidden parameter dimension of 44 (capturing variations in damping, mass, inertia, and friction). The plot reports the average per-sample $\ell_{1}$ prediction error between the true and predicted next states on the test set.
(a) Training performance Refer to caption
(b) OOD generalization performance
Figure 7: Comparison of our function encoder (FE) representation to a naive Transformer-based (TF) representation when both are used with the proposed shielding framework (all other hyperparameters and conditions remain the same). (a) In-distribution training (return vs. cost rate). (b) Out-of-distribution (OOD) evaluation (return vs. cost rate). In both settings, our method attains higher returns and fewer constraint violations. This improvement arises from the combination of a more informative latent representation and higher next-state prediction accuracy, both of which strengthen the effectiveness of the shielding mechanism.
Appendix I Additional Experiments: Safe Velocity Control in HalfCheetah
Figure 8: HalfCheetah-Velocity environment.
In this section, we evaluate whether our mechanism generalizes to different robot morphologies and task type by conducting shielding experiments in the HalfCheetah-Velocity task. We follow the safetyconstrained velocity environment from Safety Gymnasium (ji2023safety), but strengthen the safety requirement by tightening the velocity limit from 3.0 to 2.0. As in our earlier settings, the hidden environment parameters vary each episode. To induce richer dynamics variability, we modify the HalfCheetah default parameters, including friction, body segment lengths, and gear ratios with total $14$ different hidden parameters:
$\{friction,torso\_length,bthigh\_length,\cdots,foot\_gear\},$
and resample them at the beginning of each episode. During training, each parameter is scaled uniformly within $[0.7,1.3]$. For out-of-distribution (OOD) evaluation, we use the disjoint ranges $0.4,0.7]\cup[1.3,1.6]$.
We use 5 neural basis functions in the function encoder, resulting in a 5-dimensional representation that serves as a proxy $b_{\phi}$ for the hidden parameters $\phi$. As a baseline, RCPO is run with oracle access to the true hidden parameters $\phi$, whereas our method uses the same RCPO implementation without oracle information and instead augment the state with the learned basis coefficients $b_{\phi}$ as its parameter estimate. All models share the same RL hyperparameters. For hyperparameters for shielding, we use 10 samples of actions, and a safety bonus $\alpha=1$.
Across both training and OOD settings, Figure 9 shows consistent improvements in safety and overall performance. Our method achieves a favorable reward-cost Pareto frontier compared to the oracle informed baseline, demonstrating that the learned representation and shielding mechanism transfer effectively to more complex robot dynamics.
Figure 9: (Left) Return vs. cost rate during the last 20 training epochs, showing the reward–safety tradeoff achieved by each method. (Right) Out-of-distribution (OOD) evaluation performance under shifted hidden parameters.
Appendix J Why Not $Q_{C}$, But $Q_{\text{safe}}$?